Sophie is twice as old as her brother Simon and half as old as her father.
It will be another 22 years before Simon is half as old as their father.
How old is Sophie now?

Answer by lwsshak3(11628)

Sophie is twice as old as her brother Simon and half as old as her father.
It will be another 22 years before Simon is half as old as their father.
How old is Sophie now?
***
let x=Simon's age
2x=Sophie's age
4x=father's age
..
In 22 years:
x+22=Simon's age
4x+22=Father's age
4x+22=2(x+22)
4x+22=2x+44
2x=22
x=11
2x=22
Sophie's age now=22